OpenStudy (anonymous):

is 98 a perfect square ? explain why

OpenStudy (mathteacher1729):

A perfect square is a number which can be written as the product of two identical numbers: 16, 25, and 36 are prefect squares because: 4*4 = 16 5*5 = 25 and 6* 6 = 36 With that in mind, you should be able to continue multiplying numbers until you find your answer. :)

OpenStudy (anonymous):

no because \[\sqrt{98}=\]9.899

OpenStudy (anonymous):

im so confused

OpenStudy (anonymous):

????????????

OpenStudy (anonymous):

it is not a perfect square

OpenStudy (anonymous):

why is that ?

OpenStudy (anonymous):

for a number to be a perfect sqaure , the number must have a positive integer square root

OpenStudy (mathteacher1729):

Riah... finish these multiplications: 7*7 = 8*8 = 9*9 = 10* 10 = 11*11 = then ask yourself "Is 98 one of the answers?" If YES = then 98 is a "perfect square" Otherwise, it is not a perfect square. :)

OpenStudy (anonymous):

Ok i get it :)

OpenStudy (anonymous):

\[7\sqrt{2}\]

OpenStudy (anonymous):

wat ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

not a perfect square because it is 9.899 if you square root it